# Mirror partner for a Klein quartic polynomial

### v2

@@ 3.3.3. Consider $g=\mathrm{id}$
While computing the vector space ${\mathsf{HH}}^{*}(f,G)$ one has to find the $G$–invariants of $\mathrm{Jac}(f)\xi_{\mathrm{id}}$. We have $h^{\ast}(\xi_{\mathrm{id}})=\xi_{\mathrm{id}}$ and it remains to compute the $G$–invariants of $\mathrm{Jac}(f)$ itself. This goal can be achieved by symmetrization procedure. Namely, if $\{\phi_{k}({\bf x})\}_{k=1}^{\mu}$ are the monomials whose classes generate $\mathrm{Jac}(f)$, then the classes of $\{\sum_{g\in G}\phi_{k}(g\cdot{\bf x})\}_{k=1}^{\mu}$ generate $(\mathrm{Jac}(f))^{G}$. However, some of these classes might be zero or linearly dependant.
The basis monomials of $\phi_{\bullet}$ can be taken to be $x_{1}^{a_{1}}x_{2}^{a_{2}}x_{3}^{a_{3}}$, s.t. $\sum_{k}a_{k}\leq 6$ and $0\leq a_{k}\leq 2$. Also $[\mathrm{hess}(f)]=756[(x_{1}x_{2}x_{3})^{2}]$. Assign the grading $\sum_{k}a_{k}$ to each of these monomials. The vector space $\mathrm{Jac}(f)$ decomposes into the direct sum of the graded pieces. The dimensions of these pieces can be composed into the following vector
1, 3, 6, 7, 6, 3, 1.13676311,\ 3,\ 6,\ 7,\ 6,\ 3,\ 1.1 , 3 , 6 , 7 , 6 , 3 , 1 .
This vector is symmetric by its middle point — the $4$–th component because the pairing $\eta_{f}$ is non–degenerate.
@@ 4. Admissible symmetry groups
This section gathers the computations of ${\mathsf{HH}}^{*}(f,G)$ for a Klein quartic polynomial $f$ and all admissible symmetry groups $G\subset{\mathrm{GL}}_{f}$. Namely, the groups $G$, s.t. $\det g=\pm 1$ for any $g\in G$.
@@ 4.1. The subgroups
For any group $G$ from the list above we also consider its extension by ${\mathbb{Z}}/2{\mathbb{Z}}$ obtained by adding an element $-\mathrm{id}\in{\mathrm{GL}}(3,{\mathbb{C}})$. Denote such group by $\widehat{G}$. Obviously, if $g_{1},\dots,g_{r}$ are all the representatives of the conjugacy classes of $G$, then ±$\pm g_{1},\dots,\pm g_{r}$ are the representatives of the conjugacy classes of $\widehat{G}$.
In the following computations we make use of different statements of Section 3 without quoting them each time. In particular, we make use of Eq. (7) to find ${\mathsf{HH}}^{*}(f,G)$ and Section 3.2,3.3 to compute the $G$–invariants. In particular, for any $G$ beneath we find the basis of ${\mathsf{HH}}^{*}(f,G)$.
@@ 4.2. The group ${\mathrm{SL}}_{f}$
The elements $S$ and $S^{3}$ don’t have $1$ as an eigenvalue, while all other do. The multiples of the centralizers are ±$\pm 1$ for $R$, and all equal to $1$ for $T$ and $RS^{3}$. This gives
@@
Both classes are $G$–invariant and non–zero in $\mathrm{Jac}(f)$.
@@ 4.4. The group (b)
The identity sector basis is given by $11$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with $\phi(x)$ taken from the list below
@@ 4.5. The group (c)
The identity sector basis is given by $9$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 1, 2, 1, 2, 1, 1.11212111,\ 1,\ 2,\ 1,\ 2,\ 1,\ 1.1 , 1 , 2 , 1 , 2 , 1 , 1 .
@@ 4.6. The group (d)
The identity sector basis is given by $15$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 1, 4, 3, 4, 1, 1.11434111,\ 1,\ 4,\ 3,\ 4,\ 1,\ 1.1 , 1 , 4 , 3 , 4 , 1 , 1 .
@@ 4.7. The group (e) - Klein 4-group
Let $G:=\{\mathrm{id},RS^{2}RS,SRS^{6},S^{2}RS^{3}RS\}$. It’s not hard to see that $G$ is indeed homomorphic to a Klein four-group. All nontrivial group elements have eigenvalues $(1,-1,-1)$.
We have ${\mathcal{C}}^{G}=G$ and the centralizers being all $G$. For the nonidentical elements $g$, the multiples of the centralizers are ±$\pm 1$ and therefore
@@
The identity sector basis is given by $9$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 3, 1, 3, 0, 1.1 0313 011,\ 0,\ 3,\ 1,\ 3,\ 0,\ 1.1 , 0 , 3 , 1 , 3 , 0 , 1 .
@@ 4.8. The group (f) - $D_{6}$
We have ${\mathcal{C}}^{G}=\{\mathrm{id},a,b\}$. All these elements have $1$ as en eigenvalue. The corresponding centralizers are $Z(a)=\langle a\rangle$, $Z(b)=\langle b\rangle$.
@@
The identity sector basis is given by $7$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 2, 1, 2, 0, 1.1 0212 011,\ 0,\ 2,\ 1,\ 2,\ 0,\ 1.1 , 0 , 2 , 1 , 2 , 0 , 1 .
@@ 4.9. The group (g) - $D_{8}$
The multiples of the centralizers are ±$\pm 1$ for $a^{2}$, $b$ and $ab$. The multiples of the centralizers for $a$ are ±$\pm\sqrt{-1},1$. Due to this we have
@@
The identity sector basis is given by $6$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 2, 0, 2, 0, 1.1 02 02 011,\ 0,\ 2,\ 0,\ 2,\ 0,\ 1.1 , 0 , 2 , 0 , 2 , 0 , 1 .
@@ 4.10. The group (h)
The identity sector basis is given by $3$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 0, 1, 0, 0, 1.1 0 01 0 011,\ 0,\ 0,\ 1,\ 0,\ 0,\ 1.1 , 0 , 0 , 1 , 0 , 0 , 1 .
@@ 4.11. The group (i)
We have ${\mathcal{C}}^{G}=\{\mathrm{id},c_{2},c_{3},c_{4},v_{4}\}$. All the elements listed have $1$ as an eigenvalue. The centralizers are $Z(c_{2})=\{\mathrm{id},RSRS^{2},RS^{4}RS^{4},RS^{5}RS^{6}\}$ giving the multiples ±$\pm 1$; $Z(c_{3})=\{\mathrm{id},TS^{4},T^{2}S^{5}\}$ giving the multiples $1$; $Z(c_{4})=\{\mathrm{id},TRSRS^{5},TRS^{2}RS^{3},RS^{6}RS\}$ giving the multiples $1$; $Z(v_{4})=\{\mathrm{id},RS^{4}RS^{3},R,T^{2}RS^{2}RS^{6},T^{2}RS^{3}RS,T^{2}RS^{5}RS^{2},T^{2}RS^{6}RS^{4},TRSRS^{5}\}$ giving the multiples ±$\pm 1$. We have
@@
The identity sector basis is given by $4$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 1, 0, 1, 0, 1.1 01 01 011,\ 0,\ 1,\ 0,\ 1,\ 0,\ 1.1 , 0 , 1 , 0 , 1 , 0 , 1 .
@@ 4.12. The group (j)
We have ${\mathcal{C}}^{G}=\{\mathrm{id},c_{3},v_{4}\}$. All the elements listed have $1$ as an eigenvalue. The centralizers are $Z(v_{4})=\{\mathrm{id},T^{2}RS^{3}RS,T^{2}RS^{5}RS^{2},T^{2}RS^{6}RS^{4}\}$ giving the multiples ±$\pm 1$; $Z(c_{3})=\{\mathrm{id},TS^{4},T^{2}S^{5}\}$ giving the multiple $1$ exclusively. We have
@@
The identity sector basis is given by $10$ elements $[\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 3, 2, 3, 0, 1.1 0323 011,\ 0,\ 3,\ 2,\ 3,\ 0,\ 1.1 , 0 , 3 , 2 , 3 , 0 , 1 .
@@ 4.13. ${\mathbb{Z}}/2{\mathbb{Z}}$–extension of the Klein 4-group
In this section denote $\widehat{G}:=\{\pm g\ |\ g\in G\}$ with $g\in G:=\{\mathrm{id},RS^{2}RS,SRS^{6},S^{2}RS^{3}RS\}$.
Then for any $g\in G$ the eigenvalues of $g$ are $(1,-1,-1)$ and those of $-g$ are $(-1,1,1)$. We get $\mathrm{Fix}(-g)\cong{\mathbb{C}}^{2}$, however the sector ${\mathcal{B}}_{-g}=0$ as explained in Section 3.3.2. Similarly ${\mathcal{B}}_{g}=({\mathbb{C}}\langle[\widehat{x}]\rangle\xi_{g})^{\widehat{G}}=0$ because $[\widehat{x}]\xi_{g}$ is not invariant under the action of $-\mathrm{id}$.
@@
It is $8$–dimensional with the product restricted from $\mathrm{Jac}(f)$. The only non–zero products that do not involve the element $1\xi_{\mathrm{id}}$ are
@@
4.14. ${\mathbb{Z}}/2{\mathbb{Z}}$–extension of ${\mathrm{SL}}_{f}$
We have ±${\mathcal{B}}_{\pm S}={\mathcal{B}}_{\pm S^{3}}={\mathcal{B}}_{-\mathrm{id}}=0$ because the fixed locus of these elements is 00 and the respective generator $\xi_{g}$ is not preserved by the action of $-\mathrm{id}$.
The fixed loci of $R$ and $-R$ are $1$ and $2$–dimensional respectively.
and ${\mathcal{B}}_{-R}=0$ as explained in Section 3.3.2.
For $g\in\{\pm RS^{3},\pm T\}$ the fixed locus is $1$–dimensional giving
The identity sector is spanned by
We conclude that ${\mathsf{HH}}^{*}(f,\widehat{G})$ is of dimension $2\cdot 4+2=10$.

### v3

@@ 3.4.3. Consider $g=\mathrm{id}$
While computing the vector space ${\mathsf{HH}}^{*}(f,G)$ one has to find the GG–invariants of $\mathrm{Jac}(f)\xi_{\mathrm{id}}$. We have $h^{\ast}(\xi_{\mathrm{id}})=\xi_{\mathrm{id}}$ and it remains to compute the GG–invariants of $\mathrm{Jac}(f)$ itself. This goal can be achieved by symmetrization procedure. Namely, if $\{\phi_{k}({\bf x})\}_{k=1}^{\mu}$ are the monomials whose classes generate $\mathrm{Jac}(f)$, then the classes of $\{\sum_{g\in G}\phi_{k}(g\cdot{\bf x})\}_{k=1}^{\mu}$ generate $(\mathrm{Jac}(f))^{G}$. However, some of these classes might be zero or linearly dependant.
The basis monomials of $\phi_{\bullet}$ can be taken to be $x_{1}^{a_{1}}x_{2}^{a_{2}}x_{3}^{a_{3}}$, such that $\sum_{k}a_{k}\leq 6$ and $0\leq a_{k}\leq 2$. Also $\mathrm{hess}(f)]=756[(x_{1}x_{2}x_{3})^{2}]$. Assign the grading $\sum_{k}a_{k}$ to each of these monomials. The vector space $\mathrm{Jac}(f)$ decomposes into the direct sum of the graded pieces. The dimensions of these pieces can be composed into the following vector
1, 3, 6, 7, 6, 3, 1.1,\ 3,\ 6,\ 7,\ 6,\ 3,\ 1.
This vector is symmetric by its middle point — the 44–th component because the pairing $\eta_{f}$ is non–degenerate.
@@ 4. Admissible symmetry groups
This section gathers the computations of ${\mathsf{HH}}^{*}(f,G)$ for a Klein quartic polynomial $f$ and all admissible symmetry groups $G\subset{\mathrm{GL}}_{f}$. Namely, the groups $G$, such that $\det g=\pm 1$ for any $g\in G$.
@@ 4.1. The subgroups
For any group $G$ from the list above we also consider its extension by ${\mathbb{Z}}/2{\mathbb{Z}}$ obtained by adding an element $\mathrm{id}\in{\mathrm{GL}}(3,{\mathbb{C}})$. Denote such group by $\widehat{G}$. Obviously, if $g_{1},\dots,g_{r}$ are all the representatives of the conjugacy classes of $G$, then $\pm g_{1},\dots,\pm g_{r}$ are the representatives of the conjugacy classes of $\widehat{G}$.
4.2. Presentation of the results
In the following computations we make use of different statements of Section 3 without quoting them each time. In particular, we make use of Eq. (7) to find ${\mathsf{HH}}^{*}(f,G)$ and Section 3.3,3.4 to compute the GG–invariants. In particular, for any $G$ beneath we find the basis of ${\mathsf{HH}}^{*}(f,G)$.
In order to apply Eq. (7) we compute ${\mathcal{C}}^{G}$ and the centralizers $Z(g)$ for any representative $g\in{\mathcal{C}}^{G}$. Following Section 3.4.1 for $g$, such that $\det(g)=1$ we need to compute the multiples of the centralizers as in Eq. (8) for all $h\in Z(g)$. We will only list the different values of these multiples in what follows.
The identity sector is often of the highest dimension. We compute its polynomial basis every time. Note that this basis can always be chosen to be homogeneous in the polynomial degree. We provide what we call the dimension vector — the length $7$ vector consisting of the number of the degree $0,\dots,6$ basis polynomials.
@@ 4.3. The group ${\mathrm{SL}}_{f}$
The elements $S$ and $S^{3}$ don’t have $1$ as an eigenvalue, while all other do. The multiples of the centralizers are $\pm 1$ for $R$, and all equal to $1$ for $T$ and $RS^{3}$. This gives
@@
Both classes are GG–invariant and non–zero in $\mathrm{Jac}(f)$.
@@ 4.5. The group (b)
The identity sector basis is given by $11$ elements $\phi(x)]\xi_{\mathrm{id}}$ with $\phi(x)$ taken from the list below
@@ 4.6. The group (c)
The identity sector basis is given by $9$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 1, 2, 1, 2, 1, 1.1,\ 1,\ 2,\ 1,\ 2,\ 1,\ 1.
@@ 4.7. The group (d)
The identity sector basis is given by $15$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 1, 4, 3, 4, 1, 1.1,\ 1,\ 4,\ 3,\ 4,\ 1,\ 1.
@@ 4.8. The group (e) - Klein 4-group
Let $G:=\{\mathrm{id},RS^{2}RS,SRS^{6},S^{2}RS^{3}RS\}$. It’s not hard to see that $G$ is indeed homomorphic to a Klein four-group. All nontrivial group elements have eigenvalues (1,1,1)(1,-1,-1).
We have ${\mathcal{C}}^{G}=G$ and the centralizers being all $G$. For the nonidentical elements $g$, the multiples of the centralizers are $\pm 1$ and therefore
@@
The identity sector basis is given by $9$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 3, 1, 3, 0, 1.1,\ 0,\ 3,\ 1,\ 3,\ 0,\ 1.
@@ 4.9. The group (f) - $D_{6}$
We have ${\mathcal{C}}^{G}=\{\mathrm{id},a,b\}$. All these elements have $1$ as an eigenvalue. The corresponding centralizers are $Z(a)=\langle a\rangle$, $Z(b)=\langle b\rangle$.
@@
The identity sector basis is given by $7$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 2, 1, 2, 0, 1.1,\ 0,\ 2,\ 1,\ 2,\ 0,\ 1.
@@ 4.10. The group (g) - $D_{8}$
The multiples of the centralizers are $\pm 1$ for $a^{2}$, $b$ and $ab$. The multiples of the centralizers for $a$ are $\pm\sqrt{-1},1$. Due to this we have
@@
The identity sector basis is given by $6$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 2, 0, 2, 0, 1.1,\ 0,\ 2,\ 0,\ 2,\ 0,\ 1.
@@ 4.11. The group (h)
The identity sector basis is given by $3$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 0, 1, 0, 0, 1.1,\ 0,\ 0,\ 1,\ 0,\ 0,\ 1.
@@ 4.12. The group (i)
We have ${\mathcal{C}}^{G}=\{\mathrm{id},c_{2},c_{3},c_{4},v_{4}\}$. All the elements listed have $1$ as an eigenvalue. The centralizers are $Z(c_{2})=\{\mathrm{id},RSRS^{2},RS^{4}RS^{4},RS^{5}RS^{6}\}$ giving the multiples $\pm 1$; $Z(c_{3})=\{\mathrm{id},TS^{4},T^{2}S^{5}\}$ giving the multiples $1$; $Z(c_{4})=\{\mathrm{id},TRSRS^{5},TRS^{2}RS^{3},RS^{6}RS\}$ giving the multiples $1$; $Z(v_{4})=\{\mathrm{id},RS^{4}RS^{3},R,T^{2}RS^{2}RS^{6},T^{2}RS^{3}RS,T^{2}RS^{5}RS^{2},T^{2}RS^{6}RS^{4},TRSRS^{5}\}$ giving the multiples $\pm 1$. We have
@@
The identity sector basis is given by $4$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 1, 0, 1, 0, 1.1,\ 0,\ 1,\ 0,\ 1,\ 0,\ 1.
@@ 4.13. The group (j)
We have ${\mathcal{C}}^{G}=\{\mathrm{id},c_{3},v_{4}\}$. All the elements listed have $1$ as an eigenvalue. The centralizers are $Z(v_{4})=\{\mathrm{id},T^{2}RS^{3}RS,T^{2}RS^{5}RS^{2},T^{2}RS^{6}RS^{4}\}$ giving the multiples $\pm 1$; $Z(c_{3})=\{\mathrm{id},TS^{4},T^{2}S^{5}\}$ giving the multiple $1$ exclusively. We have
@@
The identity sector basis is given by $10$ elements $\phi(x)]\xi_{\mathrm{id}}$ with the dimension vector
1, 0, 3, 2, 3, 0, 1.1,\ 0,\ 3,\ 2,\ 3,\ 0,\ 1.
@@
4.14. ${\mathbb{Z}}/2{\mathbb{Z}}$–extension of ${\mathrm{SL}}_{f}$
We have ${\mathcal{B}}_{\pm S}={\mathcal{B}}_{\pm S^{3}}={\mathcal{B}}_{-\mathrm{id}}=0$ because the fixed locus of these elements is 0 and the respective generator $\xi_{g}$ is not preserved by the action of $\mathrm{id}$.
The fixed loci of $R$ and R-R are $1$ and 22–dimensional respectively.
and ${\mathcal{B}}_{-R}=0$ as explained in Section 3.4.2.
For $g\in\{\pm RS^{3},\pm T\}$ the fixed locus is 11–dimensional giving
The identity sector is spanned by
[…]
@@ 4.15. ${\mathbb{Z}}/2{\mathbb{Z}}$–extension of the Klein 4-group
In this section denote $\widehat{G}:=\{\pm g\ |\ g\in G\}$ with $G:=\{\mathrm{id},RS^{2}RS,SRS^{6},S^{2}RS^{3}RS\}$.
Then for any $g\in G$ the eigenvalues of $g$ are (1,1,1)(1,-1,-1) and those of g-g are (1,1,1)(-1,1,1). We get $\mathrm{Fix}(-g)\cong{\mathbb{C}}^{2}$, however the sector ${\mathcal{B}}_{-g}=0$ as explained in Section 3.4.2. Similarly ${\mathcal{B}}_{g}=({\mathbb{C}}\langle[\widehat{x}]\rangle\xi_{g})^{\widehat{G}}=0$ because $\widehat{x}]\xi_{g}$ is not invariant under the action of $\mathrm{id}$.
@@
[…]
@@
This suggests that the product structure for the B–model state space might be defined differenly for the algebra isomorphism to hold true.